OpenStudy (anonymous):

How to solve for n: nP5=14(nP4) and nP3=17(nP2) permutations?

Directrix (directrix):

I need to ask two questions before beginning this problem. 1) Do you understand this: 5! = 5*4! ? 2) Do you understand this: (n - 4)! = (n -4) * (n - 5)! ?

OpenStudy (anonymous):

Actually, I understand how to do the problems now. Thanks though!

Directrix (directrix):

Alrighty, then. I asked the questions because those relationships are needed to solve the problems.